OpenStudy (lilcathyx):

Medal..! The first challenge to complete is the Penny Problem. The radio station is giving the winner of this challenge a prize pack that includes tickets to see his or her favorite band in concert. To start off the challenge, the radio station has placed pennies in a cylindrical glass jar. Each penny is 0.75 inches in diameter and 0.061 inches thick. If the cylindrical glass jar containing the pennies has a diameter of 6 inches and a height of 11.5 inches, how many pennies can fit inside the jar?

OpenStudy (lilcathyx):

all i need help in is in the Pennies volume & this is what i have so far

OpenStudy (lilcathyx):

V=pi R^2h V=3.14 (0.375)^2( 0.061) is the height right ?

OpenStudy (lilcathyx):

would it be right if i multiplied 0.375 x 0.375-> (0.140625) & then make the formula as V= 3.14( 0.141)(0.061) ?

OpenStudy (lilcathyx):

@just_one_last_goodbye

OpenStudy (just_one_last_goodbye):

i cant help but i know someone that can @morganchavez

OpenStudy (lilcathyx):

Thank you :)

OpenStudy (just_one_last_goodbye):

np c:

Elsa213 (elsa213):

it can help

OpenStudy (lilcathyx):

Thanks !